OpenStudy (anonymous):

The function that has the same shape as y=x^2 but is shifted 2 units left and 1 unit up has the equation

jimthompson5910 (jim_thompson5910):

y=x^2 shifted 2 units left turns into y = (x+2)^2 shifting that 1 unit up gives you y = (x+2)^2 + 1

OpenStudy (anonymous):

how would i solve that

jimthompson5910 (jim_thompson5910):

not sure what you mean

OpenStudy (anonymous):

y= (x+2)^2 + 1 how do you solve this problem

jimthompson5910 (jim_thompson5910):

you just want the equation right?

OpenStudy (anonymous):

yeah but if i had to solve how would i solve

OpenStudy (anonymous):

this a multiply choice question but the answer you gave isnt one of the chooses

jimthompson5910 (jim_thompson5910):

what answer choices does it give you

OpenStudy (anonymous):

y=x^2+3x+5 y=x^2+4x+4 y=x^2+5x+4 y=x^2-5x+4 y=x^2+4x+5

jimthompson5910 (jim_thompson5910):

y = (x+2)^2 + 1 y = x^2+4x+4 + 1 Combine like terms

OpenStudy (anonymous):

x^2 +4x+5

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

k , thanks you are such a great help , you wanna help me a lil more

jimthompson5910 (jim_thompson5910):

yw, sure one more